The table shows the results of a survey of students in two math classes.
Find P(more than 1 hour of TV | 6th period class). Round to the nearest thousandth.
Did You Watch More Than One Hour of TV Last Night?

3rd Period Class: Yes- 11 No- 6
6th Period Class: Yes- 12 No- 10
* 0.647
* 0.565
* 0.435
* 0.765

Answer by stanbon(75887)

The table shows the results of a survey of students in two math classes.
Find P(more than 1 hour of TV | 6th period class). Round to the nearest thousandth.
-----------------------
# in 6th grade class = 22
# in 6th that watched > 1 hr. of TV = 12
----
P(> 1 | 6th) = 12/22 = 6/11 = 0.5455
-----------------------------------------
Cheers,
Stan H.
=========================================

Did You Watch More Than One Hour of TV Last Night?
3rd Period Class: Yes- 11 No- 6
6th Period Class: Yes- 12 No- 10
* 0.647
* 0.565
* 0.435
* 0.765